# A Proxy Attack-Free Strategy for Practically Improving the Poisoning Efficiency in Backdoor Attacks

### v2

@@ V-B Experiments on CIFAR-10 Dataset
Results on different data transforms and architectures. Table I provides a comprehensive overview of the attack success rates for BadNets, Blended, and Optimized triggers, considering various data transforms and architectures. Constructing the proxy attack for FUS, we employ VGG16 with RandomCrop+RandomHorizontalFlip and SGD-0.01 as the optimizer. Our proposed PFS method significantly enhances the poisoning efficacy in backdoor attacks. Across all 54 configurations, the attack success rate of poisoning samples selected through our approach consistently surpasses that of randomly selected samples at identical poisoning rates. On average, our method achieves a improvement of 0.044, 0.074, and 0.066 in terms of attack success rate for BadNets, Blended, and Optimized triggers, respectively. In general, our method outperforms FUS, demonstrating greater ASR in 48 of the 54 scenarios. Notably, FUS searches necessitate significantly more time compared to our method, often by hundreds or even thousands of times. Furthermore, when the model is VGG16 and the trigger is set to Blended, FUS surpasses our PFS. This can be attributed to two factors: (i) Both the proxy attack model employed in FUS and the model utilized in the victim phase are VGG16; (ii) In comparison to BadNets and Optimized triggers, the Blended trigger increases the similarity distance between benign and corresponding poisoning samples, thereby marginally impacting the performance of our proposed PFS. The combination of FUS and PFS, denoted as FUS+PFS, excels in 51 out of the 54 scenarios. This outcome stems from PFS effectively eliminating inefficient samples, consequently decreasing the search space for FUS.
TABLE II: The attack success rate under different optimizers on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets, Blended, and Optimized are $1\%$, $1\%$, and $0.5\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.
Trigger Optimizer Model
VGG16 ResNet18 PreAct-ResNet18
Random PFS FUS FUS+PFS Random PFS FUS FUS+PFS Random PFS FUS FUS+PFS
Results on different optimizer and architectures. Table II presents the ASR for the BadNets, Blended, and Optimized triggers, considering different optimizers and architectures. We utilize VGG16 with RandomCrop+RandomHorizontalFlip data transformations and SGD-0.01 as the optimizer for constructing the proxy attack in FUS. The outcomes underscore the effectiveness of our PFS method. Firstly, across all 36 configurations, our PFS consistently achieves higher attack success rates than random selection, with average improvements of 0.0437, 0.052 and 0.017 for the BadNets, Blended, and Optimized triggers, respectively. Secondly, our PFS outperforms FUS in 25 out of the 36 cases. Similar to previous experiments, when the model is VGG16 and the trigger is set to Blended, FUS outperforms our PFS. Notably, the combination of our method with FUS, referred to as FUS+PFS, achieves the best performance in 33 out of the 36 settings.
The Supplementary Materials contain additional experiments aimed at further evaluating the effectiveness and robustness of our proposed strategy. Specifically, in Sec. B and Sec. C, we provide the attack success rate (ASR) and benign accuracy (BA) under poisoning rates of $1.5\%$ and $2\%$ on the CIFAR-10 dataset, respectively. The outcomes illustrate the robustness of our method across different poisoning rates while maintaining benign accuracy. Moreover, in Sec. D, we present the ASR for an additional target class (class 3) on the CIFAR-10 dataset. Overall, our results consistently affirm the effectiveness of the proposed pfs and support our hypothesis that the efficiency of backdoor attacks depends on the interaction between individual similarity and set diversity.
V-C Experiments on Tiny-ImageNet Dataset
Results on different data transforms and architectures. Table III shows that PFS outperforms the random method in terms of attack success rates of poisoning samples on Tiny-ImageNet, with average boosts of 0.054, 0.094, and 0.072 on BadNets, Blended, and Optimized triggers, respectively. In comparison to FUS, PFS outperforms FUS in 45 out of 54 cases. In contrast to experiments on the CIFAR-10 dataset, our proposed PFS method consistently achieves higher average ASR than FUS across different data transformations on the Tiny-ImageNet dataset. When the two methods are combined, the best results are achieved in 34 out of 54 settings, indicating the superiority of our method. These findings provide strong evidence supporting the effectiveness of our proposed method on the scenarios involving extensive datasets.
@@ V-C Experiments on Tiny-ImageNet Dataset
Blended WaNet ISSBA
RSS 0.864 0.813 0.764
FUS [50] 0.929 0.877 0.838
OPS [17] 0.882 0.843 0.792
Gradient[10] 0.901 0.851 0.808
RD Score[49] 0.913 0.847 0.806
PFS (Ours) 0.948 0.893 0.861
@@
Dataset PFS (w/o pre-training) PFS (w pre-training) FUS
CIFAR-10 17 562 9,150
Tiny-ImageNet 19 6784 106284
@@ V-D Comparison with Recent Sample Selection Strategies
Table IV presents a comparative analysis between our PFS and recent sample selection strategies (FUS [50], OPS [17], Gradient[10], and RD Score[49]) on the Tiny-ImageNet dataset. Various recent triggers (Blended, WaNet, and ISSBA) are considered in this experiment. The results demonstrate that all sample selection strategies exhibit superior ASR compared to the random selection strategy (RSS), with our proposed PFS consistently demonstrating the highest sample efficiency across all settings when compared to other recent sample selection strategies.
@@ V-F Analysis of Time Consumption
Table V presents a comparison of the running times of our PFS with those of FUS. PFS demonstrates remarkable speed, being 538×538$\times$538 × faster than FUS on CIFAR-10 dataset. Even with pre-training time included, our method is still approximately 16×16$\times$16 × faster on CIFAR-10 dataset. Importantly, our method utilizes the same pre-trained model for different triggers, while FUS needs to conduct a complete search from the beginning for each trigger. This advantage makes our method a practical choice for situations where efficiency is crucial.
@@ V-G Ablation Study
We conduct an ablation study on the hyperparameter $m$, which determines the diversity of the poisoning sample set. The results are shown in Figure 9, and the attack success rates demonstrate similar trends under different data transforms. Initially, as diversity gradually increases, the attack success rates also increase. However, too large of a value for $m$ can reduce the role of similarity and weaken the attack. These findings highlight the importance of balancing similarity and diversity for efficient poisoning samples, supporting our claim in Sec. IV-A. In our PFS, we set $m=10$ in all experimental settings.
@@
We evaluate the robustness of PFS against backdoor defense using the pruning-based defense method [32] and the tune-based defense method FST [36]. Pruning [32] is a widely used defense mechanism. It achieves this by pruning neurons that remain inactive in response to benign inputs, thereby neutralizing the effectiveness of the backdoor behavior. On the other hand, FST [36] is a tuning-based backdoor purification method. It operates by inducing feature shifts through deliberate adjustments to the classifier weights, effectively diverging them from the initially compromised weights. As depicted in Figure 10, PFS demonstrates superior robustness against Pruning [32] compared to RSS and FUS. Similarly, Figure 11 illustrates that our proposed PFS effectively against the tune-based defense method FST [36] compared to RSS and FUS. It is noteworthy that while FST [36] demonstrates superior defense capabilities compared to Pruning [32], it comes at the expense of a substantial reduction in benign accuracy.
@@
VI Conclusion and Future Work
@@
Future Work. While the proposed PFS has demonstrated efficient data selection in backdoor attacks, we note that similarity between benign and corresponding poisoned samples does not necessarily guarantee efficiency, as similarity only filters out non-effective samples. In the future, it would be desirable to identify an unified indicator that positively correlates with data efficiency in backdoor attacks. Moreover, our proposed PFS is customized specifically for dirty-label backdoor attacks. In future research, we intend to explore efficient sample selection strategies designed specifically for clean-label backdoor attacks.
@@ Acknowledgments
The work was supported in part by the National Natural Science Foundation of China under Grands U19B2044 and 61836011.
@@ Appendix A Supplementary Materials
This supplementary material is organized as follows. We first discuss the attack success rate and benign accuracy under poisoning rates with $1.5\%$ and $2\%$ on the CIFAR-10 dataset in Sec. A-B and Sec. A-C, respectively. Additionally, we present the results of attack success rate under additional target classes on the CIFAR-10 dataset in Sec. A-D. Besides, Sec. V-C presents the experimental results on the Tiny-ImageNet dataset.
A-A The Proof of Theorem 1
@@
A-B Attack Success Rate Under Additional poisoning rates on the CIFAR-10 Dataset
We evaluate the attack success rate (ASR) under different poisoning rates on the CIFAR-10 dataset. Tables VI and VII present the ASR with trigger BadNets and Blended at poisoning rates of $1.5\%$ and $2\%$, respectively. Results under poisoning rates of $1.5\%$ and $2\%$ are similar to those at $1\%$.
@@ A-B Attack Success Rate Under Additional poisoning rates on the CIFAR-10 Dataset
TABLE VI: The attack success rate (ASR) on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $1.5\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.
@@
TABLE VII: The attack success rate (ASR) on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $2\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.
@@
A-C Benign Accuracy Under Additional poisoning rates on the CIFAR-10 Dataset
We also present the benign accuracy (BA) (i.e., the probability of classifying a benign test data to the correct label) under other poisoning rates on the CIFAR-10 dataset. Table VIII and Table IX illustrate the benign accuracy with trigger BadNets and Blended under poisoning rates of $1.5\%$ and $2\%$, respectively. The results show the similar BA with four strategies, which demonstrates that our proposed strategy has no negative impact on benign accuracy.
TABLE VIII: The benign accuracy (BA) on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $1.5\%$, respectively.
@@ A-C Benign Accuracy Under Additional poisoning rates on the CIFAR-10 Dataset
TABLE IX: The benign accuracy (BA) on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $2\%$, respectively.
@@
A-D Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
This section presents the attack success rate (ASR) under additional target classes on the CIFAR-10 dataset. Table X illustrates the attack success rate with trigger BadNets and Blended under target category 3. Compared with random selection, our PFS approach outperforms in all cases, with average boosts of 0.045 and 0.076 on BadNets and Blended triggers. The PFS method also achieves better attack results in most cases compared to FUS (22/24). The combination of these two achieves superior results in all 24 settings. These outcomes consistently confirm the conclusions we reached in the paper.
TABLE X: The attack success rate (ASR) on the CIFAR-10 dataset. In this experiment, attack target k is set to category 3. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $1\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.

### v3

@@ V-B Experiments on CIFAR-10 Dataset
Results on different data transforms and architectures. Table II provides a comprehensive overview of the ASR for BadNets, Blended, ISSBA, and Optimized triggers, considering various data transforms and architectures. Constructing the proxy attack for FUS, we employ VGG16 with RandomCrop+RandomHorizontalFlip and SGD-0.01 as the optimizer. Our proposed PFS method significantly enhances the poisoning efficacy in backdoor attacks. Across all 72 configurations, the ASR of poisoning samples selected through our approach consistently surpasses that of randomly selected samples at identical poisoning rates. On average, our method achieves a improvement of 0.044, 0.074, 0.08, and 0.066 in terms of ASR for BadNets, Blended, ISSBA, and Optimized triggers, respectively. In general, our method outperforms FUS, demonstrating greater ASR in 63 of the 72 scenarios. Notably, FUS searches necessitate significantly more time compared to our method, often by hundreds or even thousands of times. Furthermore, when the model is VGG16 and the trigger is set to Blended, FUS surpasses our PFS. This can be attributed to two factors: (i) Both the proxy attack model employed in FUS and the model utilized in the victim phase are VGG16; (ii) In comparison to BadNets and Optimized triggers, the Blended trigger increases the similarity distance between benign and corresponding poisoning samples, thereby marginally impacting the performance of our proposed PFS. The combination of FUS and PFS, denoted as FUS+PFS, excels in 69 out of the 72 scenarios. This outcome stems from PFS effectively eliminating inefficient samples, consequently decreasing the search space for FUS.
@@
Blended WaNet ISSBA BATT
RSS 0.864 0.813 0.764 0.852
FUS [63] 0.929 0.877 0.838 0.901
OPS [21] 0.882 0.843 0.792 0.869
Gradient[13] 0.901 0.851 0.808 0.880
RD Score[62] 0.913 0.847 0.806 0.893
PFS (Ours) 0.948 0.893 0.861 0.935
Results on different optimizer and architectures. Table I presents the ASR for the BadNets, Blended, ISSBA, and Optimized triggers, considering different optimizers and architectures. We utilize VGG16 with RandomCrop+RandomHorizontalFlip data transformations and SGD-0.01 as the optimizer for constructing the proxy attack in FUS. The outcomes underscore the effectiveness of PFS method. Firstly, across all 48 configurations, PFS consistently achieves higher ASR than random selection, with average improvements of 0.0437, 0.052, 0.065, and 0.017 for the BadNets, Blended, ISSBA and Optimized triggers, respectively. Secondly, PFS outperforms FUS in 36 out of the 48 cases. Similar to previous experiments, when the model is VGG16 and the trigger is set to Blended, FUS outperforms PFS. Notably, the combination of our method with FUS, referred to as FUS+PFS, achieves the best performance in 45 out of the 48 settings.
The Supplementary Materials contain additional experiments aimed at further evaluating the effectiveness and robustness of our proposed strategy. Specifically, in Sec. C and Sec. D, we provide the ASR and benign accuracy (BA) under poisoning rates of $1.5\%$ and $2\%$ on the CIFAR-10 dataset, respectively. The outcomes illustrate the robustness of our method across different poisoning rates while maintaining benign accuracy. Moreover, in Sec. E, we present the ASR for an additional target class (class 3) on the CIFAR-10 dataset. Overall, our results consistently affirm the effectiveness of the proposed pfs and support our hypothesis that the efficiency of backdoor attacks depends on the interaction between individual similarity and set diversity.
V-C Experiments on Tiny-ImageNet Dataset
Results on different data transforms and architectures. Table III shows that PFS outperforms the random method in terms of ASR of poisoning samples on Tiny-ImageNet, with average boosts of 0.054, 0.094, and 0.072 on BadNets, Blended, and Optimized triggers, respectively. In comparison to FUS, PFS outperforms FUS in 45 out of 54 cases. In contrast to experiments on the CIFAR-10 dataset, our proposed PFS method consistently achieves higher average ASR than FUS across different data transformations on the Tiny-ImageNet dataset. When the two methods are combined, the best results are achieved in 34 out of 54 settings, indicating the superiority of our method. These findings provide strong evidence supporting the effectiveness of our proposed method on the scenarios involving extensive datasets.
@@ V-C Experiments on Tiny-ImageNet Dataset
Dataset CIFAR-10 Tiny-ImageNet
PFS (w/o pre-training) 17 19
PFS (w pre-training) 562 6784
FUS [63] 9150 106284
Gradient [13] 623 7395
OPS [21] 635 7487
@@ V-D Comparison with Recent Sample Selection Strategies
Table IV presents a comparative analysis between our PFS and recent sample selection strategies (FUS [63], OPS [21], Gradient[13], and RD Score[62]) on the Tiny-ImageNet dataset. Various recent triggers (Blended, WaNet, ISSBA, and BATT) are considered in this experiment. The results demonstrate that all sample selection strategies exhibit superior ASR compared to the random selection strategy (RSS), with our proposed PFS consistently demonstrating the highest sample efficiency across all settings when compared to other recent sample selection strategies.
@@ V-F Analysis of Time Consumption
Table V presents a comparison of the running times of our PFS with those of FUS, Gradient, and OPS. PFS demonstrates remarkable speed, being 538×538$\times$538 × faster than FUS on CIFAR-10 dataset. Even with pre-training time included, our method is still approximately 16×16$\times$16 × faster on CIFAR-10 dataset. Importantly, our method utilizes the same pre-trained model for different triggers, while FUS needs to conduct a complete search from the beginning for each trigger. This advantage makes our method a practical choice for situations where efficiency is crucial.
@@ V-G Ablation Study
We conduct an ablation study on the hyperparameter $m$, which determines the diversity of the poisoning sample set. The results are shown in Figure 8, and the ASR demonstrate similar trends under different data transforms. Initially, as diversity gradually increases, the ASR also increase. However, too large of a value for $m$ can reduce the role of similarity and weaken the attack. These findings highlight the importance of balancing similarity and diversity for efficient poisoning samples, supporting our claim in Sec. III-B. In our PFS, we set $m=10$ in all experimental settings.
@@
We evaluate the robustness of PFS against backdoor defense using the pruning-based defense method [44], the tune-based defense method FST [48], and input-level backdoor detection SCALE-UP [19]. Pruning [44] is a widely used defense mechanism. It achieves this by pruning neurons that remain inactive in response to benign inputs, thereby neutralizing the effectiveness of the backdoor behavior. FST [48] is a tuning-based backdoor purification method. It operates by inducing feature shifts through deliberate adjustments to the classifier weights, effectively diverging them from the initially compromised weights. SCALE-UP [19] only requires the predicted labels to realize the effective black-box input-level backdoor detection. Specifically, it identifies and filters malicious testing samples by analyzing their prediction consistency during the pixel-wise amplification process. Furthermore, defender adopts the filtered benign samples to train the model. As depicted in Figure 9, PFS demonstrates superior robustness against Pruning [44] compared to RSS and FUS. Similarly, Figure 10 illustrates that our proposed PFS effectively against the tune-based defense method FST [48] compared to RSS and FUS. It is noteworthy that while FST [48] demonstrates superior defense capabilities compared to Pruning [44], it comes at the expense of a substantial reduction in benign accuracy. Additionally, we adopt the data-free scaled prediction consistency analysis to evaluate the effectiveness of different attack methods against the SCALE-UP defense method [19]. As shown in Table VI, our proposed PFS method demonstrates superior performance in countering the backdoor detection method SCALE-UP compared to other selection baselines.
@@
TABLE VI: Defense results (ASR) of SCALE-UP on the Tiny ImageNet dataset. The poisoning ratio is set to 1%.
Attack BadNets Blended WaNet ISSBA
RSS 0.327 0.592 0.285 0.233
Gradient 0.363 0.633 0.348 0.294
FUS 0.411 0.661 0.384 0.316
PFS 0.452 0.685 0.415 0.337
VI Conclusion and Limitations
@@
Limitations and Future Work. (i) While the proposed PFS has demonstrated efficient data selection in backdoor attacks, we note that similarity between benign and corresponding poisoned samples does not necessarily guarantee efficiency, as similarity only filters out non-effective samples. In the future, it would be desirable to identify an unified indicator that positively correlates with data efficiency in backdoor attacks. (ii) Moreover, our proposed PFS is customized specifically for dirty-label backdoor attacks. In future research, we intend to explore efficient sample selection strategies designed specifically for clean-label backdoor attacks. (iii) Our proposed PFS does not show significant improvement over the FUS method against backdoor defenses. Therefore, we will investigate more efficient sample selection methods to enhance efficiency against backdoor defenses in future studies.
@@ Acknowledgments
The work was supported by the National Natural Science Foundation of China under Grands U19B2044, U22B2062, 62172232, and 61836011; by the Jiangsu Basic Research Programs-Natural Science Foundation under grant numbers BK20200039; by the Collaborative Innovation Center of Atmospheric Environment and Equipment Technology fund.
@@ Appendix A Supplementary Materials
This supplementary material is organized as follows. We first provide the theoretical analyze of the proposed PFS in terms of active learning in Sec. A-A and the proof the Theorem 1 in Sec. A-B. Additionally, we discuss the attack success rate and benign accuracy under poisoning rates with $1.5\%$ and $2\%$ on the CIFAR-10 dataset in Sec. A-C and Sec. A-D, respectively. Besides, we present the results of attack success rate under additional target classes on the CIFAR-10 dataset in Sec. A-E. Finally, Sec. A-F presents the experimental results on the CIFAR-100 dataset.
A-A Perspective of Active Learning
The process of selecting efficient poisoning samples in the context of backdoor attacks can be analogized to the principles of active learning. Active learning (AL) is a machine learning paradigm that efficiently acquires annotated data from an extensive pool of unlabeled data [47]. Its objective is to prioritize human labeling efforts on the most informative data points, leading to improved model performance while reducing data annotation costs. Similarly, the task of selecting efficient poisoning samples in a backdoor attack aims to identify the most informative data points that can optimize the efficiency of the poisoning process, thereby minimizing the costs for the attacker. A pivotal aspect of active learning encompasses the concepts of uncertainty and diversity in the dataset [9]. Methods that rely on uncertainty utilize the predictive confidence of model to select challenging examples, while diversity-based sampling leverages the dissimilarity among data points, often involving clustering techniques. Notably, the combination of uncertainty and diversity strategies has demonstrated significant success in active learning [70, 54]. Similarly, our proposed PFS takes into account both uncertainty (Line 8 in Algorithm 1) and diversity (Line 9 in Algorithm 1) of data points.
A-B The Proof of Theorem 1
@@
A-C Attack Success Rate Under Additional poisoning rates on the CIFAR-10 Dataset
We evaluate the attack success rate (ASR) under different poisoning rates on the CIFAR-10 dataset. Tables VII and VIII present the ASR with trigger BadNets and Blended at poisoning rates of $1.5\%$ and $2\%$, respectively. Results under poisoning rates of $1.5\%$ and $2\%$ are similar to those at $1\%$.
@@ A-C Attack Success Rate Under Additional poisoning rates on the CIFAR-10 Dataset
TABLE VII: The attack success rate (ASR) on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $1.5\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.
@@
TABLE VIII: The attack success rate (ASR) on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $2\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.
@@
A-D Benign Accuracy Under Additional poisoning rates on the CIFAR-10 Dataset
We also present the benign accuracy (BA) (i.e., the probability of classifying a benign test data to the correct label) under other poisoning rates on the CIFAR-10 dataset. Table IX and Table X illustrate the benign accuracy with trigger BadNets and Blended under poisoning rates of $1.5\%$ and $2\%$, respectively. The results show the similar BA with four strategies, which demonstrates that our proposed strategy has no negative impact on benign accuracy.
TABLE IX: The benign accuracy (BA) on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $1.5\%$, respectively.
@@ A-D Benign Accuracy Under Additional poisoning rates on the CIFAR-10 Dataset
TABLE X: The benign accuracy (BA) on the CIFAR-10 dataset. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $2\%$, respectively.
@@
A-E Attack Success Rate Under Additional Target Classes on the CIFAR-10 Dataset
This section presents the attack success rate (ASR) under additional target classes on the CIFAR-10 dataset. Table XI illustrates the attack success rate with trigger BadNets and Blended under target category 3. Compared with random selection, our PFS approach outperforms in all cases, with average boosts of 0.045 and 0.076 on BadNets and Blended triggers. The PFS method also achieves better attack results in most cases compared to FUS (22/24). The combination of these two achieves superior results in all 24 settings. These outcomes consistently confirm the conclusions we reached in the paper.
TABLE XI: The attack success rate (ASR) on the CIFAR-10 dataset. In this experiment, attack target k is set to category 3. All results are averaged over 5 different runs. The poisoning rates of experiments with trigger BadNets and Blended are $1\%$, respectively. Among four select strategies, the best result is denoted in boldface while the underline indicates the second-best result. The settings corresponding to the gray background are represented as the proxy poisoning attack within the attacker phase and the actual poisoning process within the victim phase being consistent in FUS.